# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multi-modal Foundation Models

### v4

@@ 5.5 Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent representation along a latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. The disentangled latent variable would be invariant with respect to the variations in another latent dimension. In Figure 3, for each latent direction, we pass two image sequences and an inductive bias prompt based on the disentanglement formula to the MLLM to obtain a common latent explanation. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show "no clear explanation" or wrong explanations as they only align with one image sequence but do not reflect the common pattern in both image sequences. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables to capture the actual meaning.
@@
In Figure 11, we provide the "young appearance" prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 12, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Appendix D.
@@ 6 Conclusion
In this paper, we introduced LatentExplainer, a framework designed to generate semantically meaningful explanations of latent variables in deep generative models. Our experiments quantitatively revealed that the inclusion of inductive bias prompts significantly enhances the quality of explanations generated by LatentExplainer. In qualitative evaluations, LatentExplainer provided clear explanations for the latent variables, effectively reflecting superior performance when inductive bias prompts were utilized. The quantitative analysis further confirmed the necessity of inductive bias prompts and uncertainty quantification in producing accurate and consistent explanations.
7 Limitation
In this work, we use GPT-4o as our MLLM agent to explain the latent representations. Although GPT-4o has improved vision capabilities across most tasks, we still find it incapable of identifying intricate facial features, positions, and orientations. It is necessary to enhance visual understanding to improve the explanation quality for MLLM.
@@ Appendix A Latent Variable Perturbation
For VAE models, every time we target at each latent variable. The targeted latent variable is changed from -3 to 3, which is the 3 times the standard deviation of the standard Gaussian distribution of a latent variable, with a step size of 1.5. All the other latent variables are unchanged. For diffusion models, we follow $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is the strength. We set the value of $\gamma$ to 0.1, 0.2, 0.3, 0.4, and 0.5 for DDPM, and to 1, 2, 3, 4, and 5 for Stable Diffusion.
@@ Appendix B Uncertainty-aware Implementations
We denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
@@
By solving this equation across all datasets, we find the threshold $\varepsilon=0.2617$. Since $\varepsilon$ is solved across all datasets, it reduces the dependence on any specific dataset and can be applied to new data. Our final output $\mathcal{T}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\mathcal{T}$ will be "no clear explanation".
Appendix C An example of automatically generating prompts from an inductive bias formula
@@
Prompt: These two rows of images show the same pattern of change despite other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation".
@@
$p\left(z_{i}\mid z_{j}=\alpha\right)=p\left(z_{i}\mid z_{j}=\beta\right),\forall z_{i}\in G,z_{j}\in G^{\prime},G\neq G^{\prime},\alpha\neq\beta$.
$p\left(z_{i}\mid z_{i^{\prime}}=\alpha\right)\neq p\left(z_{i}\mid z_{i^{\prime}}=\beta\right),\forall z_{i},z_{i}^{\prime}\in G,i\neq i^{\prime},\alpha\neq\beta.$
Prompt: The pattern of change is associated with a group. The first two rows of images show the same pattern of change despite other variations in another group. The pattern of change in the last two rows of images should change given other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation".
@@
$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}$.
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$t=t_{1},t=t_{2},\forall t_{1}\neq t_{2}"$ is “different timesteps"""", “"absent"$\neq$"≠ " is “change"""". Please give a prompt in plain language.
Appendix D More Qualitative Evaluation Results
@@
Appendix E Visualization of the Generated Explanations
@@
Model Method AFHQ LSUN-Church CelebA-HQ
[…]
DDPM (Disentanglement Bias) GPT-4o Baseline 3.08 14.83 14.22 8.78 87.55 -3.39 5.52 23.06 16.85 11.23 88.47 -3.41 0.11 3.54 4.44 1.22 85.97 -3.16
+ LatentExplainer w/o IB 2.14 11.49 12.01 7.82 87.17 -3.41 8.78 27.01 20.15 14.08 88.85 -3.37 0.19 4.17 4.34 1.80 85.88 -3.16
+ LatentExplainer w/o UQ 16.12 31.57 31.74 21.70 89.99 -3.12 21.19 37.03 29.40 23.65 90.49 -3.19 13.11 27.38 25.97 14.03 89.15 -2.92
+ LatentExplainer 25.91 37.84 41.52 29.87 91.48 -2.97 30.92 44.21 36.39 29.23 91.80 -3.06 18.49 35.27 32.76 18.81 90.30 -2.90
Stable Diffusion (Conditional Bias) GPT-4o Baseline 7.61 26.32 21.52 17.19 90.04 -3.32 7.80 26.73 23.40 14.28 89.44 -3.46 5.79 23.89 20.44 12.68 89.63 -3.04
+ LatentExplainer w/o IB 9.12 26.59 22.46 16.28 89.77 -3.29 10.07 28.31 25.28 15.85 89.92 -3.42 5.87 24.06 21.09 12.17 89.15 -3.05
+ LatentExplainer w/o UQ 12.36 27.73 22.64 20.01 90.03 -3.18 17.19 35.33 30.50 21.76 90.73 -3.31 15.01 35.43 27.96 23.99 91.07 -2.87
+ LatentExplainer 13.27 29.98 23.56 20.08 90.17 -3.19 18.33 38.14 32.46 25.01 90.99 -3.23 18.50 40.85 33.90 23.90 91.75 -2.82
Table 5: Quantitative results for diffusion models across datasets. B represents BLEU@4, R represents ROUGE-L, M represents METEOR, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
Model Method 3DShapes CelebA-HQ dSprites
[…]
$\beta$-TCVAE (Disentanglement Bias) GPT-4o Baseline 5.51 29.77 19.12 7.92 89.18 -3.13 11.09 37.14 28.80 17.24 90.55 -3.07 0.00 16.19 10.38 7.04 87.17 -3.22
+ LatentExplainer w/o IB 5.41 31.66 22.28 10.07 89.44 -3.13 6.14 32.73 25.33 14.04 89.92 -3.11 0.00 18.73 14.42 7.54 87.42 -3.20
+ LatentExplainer w/o UQ 16.99 37.37 30.72 22.89 90.83 -2.80 21.95 48.84 39.70 22.60 92.09 -2.88 17.16 37.19 31.11 22.40 90.18 -2.89
+ LatentExplainer 25.40 49.06 32.72 22.78 91.90 -2.75 29.93 55.68 44.32 30.17 93.55 -2.77 12.55 37.30 29.86 21.69 90.28 -2.87
CSVAE (Combination Bias) GPT-4o Baseline 6.93 25.07 19.04 9.70 89.42 -2.87 10.44 37.42 28.51 15.06 90.53 -3.01 0.00 21.61 13.17 6.99 87.19 -3.17
+ LatentExplainer w/o IB 14.22 28.57 25.12 11.79 90.02 -2.85 11.96 37.02 28.51 17.49 90.50 -3.01 0.00 16.36 10.12 7.26 86.96 -3.24
+ LatentExplainer w/o UQ 34.55 39.62 33.42 30.30 90.83 -2.62 14.08 38.68 30.13 22.49 91.11 -2.87 0.00 28.49 19.19 15.65 89.78 -2.91
+ LatentExplainer 36.18 43.58 36.60 36.75 91.72 -2.52 25.34 45.18 39.82 26.96 92.09 -2.81 16.03 35.85 29.23 21.82 90.05 -2.90
CSVAE (Conditional Bias) GPT-4o Baseline 10.97 30.55 23.90 8.83 90.10 -3.02 8.46 28.78 21.93 15.78 89.62 -2.98 0.00 9.62 5.27 1.03 85.74 -3.34
+ LatentExplainer w/o IB 19.28 39.90 35.47 15.74 90.87 -2.88 8.35 28.75 18.45 12.11 89.60 -2.97 0.00 11.38 9.35 5.26 86.47 -3.29
+ LatentExplainer w/o UQ 16.73 32.28 25.17 16.19 89.84 -2.86 13.35 37.20 25.40 20.05 89.89 -2.92 5.39 19.20 15.65 6.36 86.94 -3.26
+ LatentExplainer 25.71 40.00 31.59 20.67 90.88 -2.79 24.21 42.88 36.19 19.61 91.21 -2.79 21.17 34.99 28.71 20.71 89.49 -3.08
Table 6: Quantitative results for VAE models across datasets. B represents BLEU@4, R represents ROUGE-L, M represents METEOR, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.

### v5

@@ 5.5 Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent representation along a latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. The disentangled latent variable would be invariant with respect to the variations in another latent dimension. In Figure 4, for each latent direction, we pass two image sequences and an inductive bias prompt based on the disentanglement formula to the MLLM to obtain a common latent explanation. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show “no clear explanation” or wrong explanations as they only align with one image sequence but do not reflect the common pattern in both image sequences in view of the inductive bias. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables to capture the actual meaning.
@@
In Figure 11, we provide the “young appearance” prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 12, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Appendix E.
Moreover, we also evaluate the effect of automatically generating inductive bias prompts from the inductive bias formulas within three aforementioned common bias types and the optional information about the symbol $\mathcal{I}$, as shown in Appendix Table 4, and the results are very promising. These examples demonstrate that our generated prompts not only can capture the nuance in the original inductive bias formulas but also help MLLMs generate more precise explanations.
@@ 6 Conclusion
In this paper, we introduced LatentExplainer, a framework designed to generate semantically meaningful explanations of latent variables in deep generative models. Our work makes three key contributions: (1) Inferring the meaning of latent variables by translating inductive bias formulas into structured perturbations of latent variables through a coding agent; (2) Aligning explanations with inductive biases by converting mathematical formulations into textual prompts in MLLMs; (3)Introducing an uncertainty-aware approach that assesses explanation consistency. Quantitative and qualitative evaluations across multiple datasets and generative models demonstrate that LatentExplainer significantly outperforms baseline methods. The incorporation of inductive bias prompts leads to more structured and meaningful explanations, while uncertainty-aware filtering further enhances consistency and reliability. Our findings highlight the importance of inductive bias prompting and uncertainty quantification in bridging the gap between generative models and human interpretability.
@@ Appendix A Latent Variable Perturbation
For VAE models, every time we target at each latent variable. The targeted latent variable is changed from -3 to 3, which is the 3 times the standard deviation of the standard Gaussian distribution of a latent variable, with a step size of 1.5. All the other latent variables are unchanged. The value of a specific latent variable among other latent variables can be determined by decomposing the inductive bias formulas. For diffusion models, we follow $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is the strength and $\mathcal{G}$ is a diffusion decoder. We set the value of $\gamma$ to 0.1, 0.2, 0.3, 0.4, and 0.5 for DDPM, and to 1, 2, 3, 4, and 5 for Stable Diffusion. $\mathbf{v}$ may be determined by decomposing the inductive bias formulas.
@@ Appendix B Uncertainty-aware Implementations
We follow Zhu et al. [2024] to denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
@@
By solving this equation across all datasets, we find the threshold $\varepsilon=0.2617$. Since $\varepsilon$ is solved across all datasets, it reduces the dependence on any specific dataset. Our final output $\hat{r}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\hat{r}$ will be “no clear explanation”.
Appendix C Human Annotaions
The ground-truth annotations of explanations are performed by four annotators from the United States and China. All annotators are students that had at least an undergraduate degree. Annotators were presented with the same images and the prompts as MLLMs and were asked to annotate the pattern of the images. If there is no clear pattern, just write “No clear explanation”. The annotations are then aggregated as references to calculate the automated evaluation metrics.
Appendix D An example of automatically generating prompts from an inductive bias formula
@@
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$p(z_{i}\mid z_{i^{\prime}}),\forall i\neq i^{\prime}"$ is “other variations"""", “= """" is “same"""".
Prompt: These two rows of images show the same pattern of change despite other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation”.
@@
[…]
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$p(z_{i}\mid z_{i^{\prime}}),\forall i\neq i^{\prime}"$ is “other variations"""", “G"""" is “a group"""", “"absent"$\in$"∈ " is “associated with"""", “= """" is “same"""", “"absent"$\neq$"≠ " is “change"""".
Prompt: The pattern of change is associated with a group. The first two rows of images show the same pattern of change despite other variations in another group. The pattern of change in the last two rows of images should change given other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation”.
@@
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$p_{k}=\alpha,p_{k}=\beta,\forall\alpha\neq\beta"$ is “other variations in the property of interests"""", “G"""" is “a group"""", “"absent"$\in$"∈ " is “associated with"""", “"absent"$\notin$"∉ " is “not associated with"""", “= """" is “same"""", “"absent"$\neq$"≠ " is “change"""".
@@
$p(z_{1}\mid z_{3}=\alpha)=p(z_{1}\mid z_{3}=\beta),\forall\alpha\neq\beta$.
$p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""",“$z_{3}=\alpha,z_{3}=\beta,\forall\alpha\neq\beta"$ is “other variations in the third latent variable"""", “="absent"="= " is “same"""". Please give a prompt in plain language.
Output:
Prompt: The pattern of change in the first latent variable remains the same, regardless of variations in the third latent variable. Describe the pattern of change in the first latent variable. If there is no clear pattern, just write “No clear explanation.”
Appendix E More Qualitative Evaluation Results
@@
Appendix F Visualization of the Generated Explanations
@@
Appendix G Case Study
More specialized cases of automatically generated inductive bias prompts from the inductive bias formulas within the three common types can be found in Table 4.
Formula Symbol-to-word Mapping Generated Prompt
$p(z_{1}\mid z_{3}=\alpha)=p(z_{1}\mid z_{3}=\beta),\forall\alpha\neq\beta$. $p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""",“$z_{3}=\alpha,z_{3}=\beta,\forall\alpha\neq\beta"$ is “other variations in the third latent variable"""", “=” is “same"""". The pattern of change in the first latent variable remains the same, regardless of variations in the third latent variable. Describe the pattern of change in the first latent variable. If there is no clear pattern, just write “No clear explanation.”
$p\left(z_{1}\mid z_{2}=\alpha\right)=p\left(z_{1}\mid z_{2}=\beta\right),\forall z_{1}\in G,z_{2}\in G^{\prime},G\neq G^{\prime},\alpha\neq\beta.$ $p\left(z_{1}\mid z_{3}=\alpha\right)\neq p\left(z_{1}\mid z_{3}=\beta\right),\newline \forall z_{1},z_{3}\in G,\alpha\neq\beta.$ $p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""", “$z_{2}=\alpha,z_{2}=\beta,\forall\alpha\neq\beta"$ is “other variations in the second latent variable"""", “$z_{3}=\alpha,z_{3}=\beta,\forall\alpha\neq\beta"$ is “other variations in the third latent variable"""", “G” is “a group”, “$G^{\prime}$” is “another group”, $\in$” is “associated with"""", $\notin$” is “not associated with"""", “=” is “same"""", $\neq$” is “change"""". The pattern of change in the first latent variable is associated with its group but not with another group. This means that the first latent variable remains the same despite variations in the second latent variable (which belongs to another group), but it changes when the third latent variable (which belongs to the same group) changes. Describe the pattern of change in the first latent variable. If there is no clear pattern, write “No clear explanation.”
$p\left(z_{1}\mid p_{k}=\alpha\right)\neq p\left(z_{1}\mid p_{k}=\beta\right),\forall\alpha\neq\beta.$
$\mathcal{I}:p_{k}=\text{age}.$
$p\left(z_{1}\mid\cdot\right)"$ is “pattern of change in the first latent variable"""", “$p_{k}=\alpha,p_{k}=\beta,\forall\alpha\neq\beta"$ is “other variations in the property of interests"""", “$p_{k}$” is “age”, "absent"$\neq$"≠ " is “change"""". The pattern of change in the first latent variable is associated with age. This means that as the property of interest (age) changes, the first latent variable also changes. Describe the pattern of change in the first latent variable. If there is no clear pattern, write “No clear explanation.”
Table 4: Examples of automatically generated inductive bias prompts